OpenStudy (anonymous):

A ball is kicked at a 50 degree angle with an initial velocity of 21 ft/s. Write the equation that models the height. What is the maximum height reached by the object?

OpenStudy (anonymous):

\[Hmax = \frac{ u^2 \sin^2 \theta }{ 2g }\]

OpenStudy (anonymous):

Using S = ut + at^2/2 Height = usin theta * t - g t^2/2

OpenStudy (anonymous):

im supposed to use this equation h(t) =-16t^2+v Sin (50)+h

OpenStudy (anonymous):

did you calculate the y component of the initial velocity

OpenStudy (anonymous):

equation:h(t)= -16t^2+16.08t+0 For the max height i got 101.8 ft

OpenStudy (anonymous):

is it starting from initial height of 0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Would that be correct?

OpenStudy (anonymous):

not sure, doesnt look right. how did you get that?

OpenStudy (anonymous):

\[h = 16.08t - 16t^{2}\] the max height occurs at time t = 0.5025 and the max height is onlly about 4.04 ft

OpenStudy (anonymous):

i used -b/2a wich would be -16/-32=0.5 and plug that back in the equation

OpenStudy (anonymous):

-b/2a gives -16.08/(2(-16)) = 0.5025

OpenStudy (anonymous):

plugging that in gives a max height of 4.04 ft ?

OpenStudy (anonymous):

oh i see where i made my mistake.

OpenStudy (anonymous):

Im supposed make up a problem what should i change to make it sound like a more realistic height?

OpenStudy (anonymous):

i mean a higher height

OpenStudy (anonymous):

change the angle to something steeper, and add more initial velocity

OpenStudy (anonymous):

how high are you trying to kick it ?

OpenStudy (anonymous):

Something like 9 9ft

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Thank you very much sir. :)